# Tendon-Based Proprioception in an Anthropomorphic Underactuated Robotic Hand with Series Elastic Actuators

### v1

@@ Abstract
Anthropomorphic underactuated hands are widely employed for their versatility and structural simplicity. In such systems, compact sensing integration and proper interpretation aligned with underactuation are crucial for realizing practical grasp functionalities. This study proposes an anthropomorphic underactuated hand that achieves comprehensive situational awareness of hand–object interaction, utilizing tendon-based proprioception provided by series elastic actuators (SEAs). We developed a compact SEA with high accuracy and reliability that can be seamlessly integrated into sensorless fingers. By coupling proprioceptive sensing with potential energy-based modeling, the system estimates key grasp-related variables, including contact timing, joint angles, relative object stiffness, and finger configuration changes indicating external disturbances. These estimated variables enable grasp posture reconstruction, safe handling of deformable objects, and blind grasping with proprioceptive-only recognition of objects with varying geometry and stiffness. Finger-level experiments and hand-level demonstrations confirmed the effectiveness of the proposed approach. The results demonstrate that tendon-based proprioception serves as a compact and robust sensing modality for practical manipulation without reliance on vision or tactile feedback.
@@ I Introduction
Anthropomorphic robotic hands have been widely adopted to replicate the functionality of the human hand. Among various actuation strategies, underactuated hands are extensively employed due to their structural simplicity, lightweight construction, and adaptability to diverse object geometries [1, 2]. As manipulation fundamentally involves interaction between the hand and the object, it is essential to sense grasp states throughout the interaction [3]. For instance, capturing finger configuration and force information remains essential for understanding object properties and interaction dynamics, thereby supporting the decision-making process during manipulation [4]. In underactuated hands, however, inherent uncertainty arising from interdependent joint motions complicates grasp state sensing [1, 5]. Accordingly, sensing approaches tailored to the distinctive features of underactuated mechanisms are required.
Human manipulation relies mainly on sensory feedback from vision, touch, and proprioception [6, 7], and these modalities have also been exploited in robotic hands. Vision provides global spatial information but limited force-related feedback and is highly sensitive to occlusion and lighting [8, 9]. To overcome these limitations of vision, robotic hands have primarily relied on touch, which provides direct contact information at the object interface [10]. However, integrating bulky tactile sensors into fingers undermines the inherent advantages of underactuated mechanisms, which lie in their compact and simple structure [11]. Moreover, since tactile sensors directly experience external forces through contact, they face challenges with durability and reliability [11, 12]. On the other hand, proprioception within the human sensorimotor system provides internal information about force and position [13]. Extending this concept to robotic hands enables sensing without external hardware, thereby facilitating simpler designs, easier assembly, and improved maintainability [14]. Furthermore, because proprioception is fundamentally decoupled from direct environmental interactions, it delivers robust performance that remains unaffected by contact conditions [15, 16].
@@
In previous work, motor current has been used as a surrogate for proprioception, but its accuracy is constrained by electrical noise and transmission friction, which obscure reliable force estimation [17, 18]. Therefore, more direct approaches–such as load cells [15] or torque sensors [5, 19], fiber Bragg gratings (FBG) sensors [14], and series elastic actuators (SEAs) [16, 20, 21, 22, 23]–have been explored for measuring internal actuation force. However, applying them to anthropomorphic underactuated hands introduces two major challenges. First, achieving compact integration of proprioceptive sensing mechanisms while ensuring high sensing performance is challenging. Existing implementations often take the form of bulky grippers or hands with enlarged bases and palms, limiting their applicability in anthropomorphic hands [16, 19, 20, 21]. Second, previous studies lack a comprehensive interpretation and utilization of the implicit force signals of underactuation, which restricts practical grasp functionalities. Some works reported force profiles for object contact or stiffness, but the results were largely illustrative with little attention paid to the adaptive characteristics of underactuation [14, 21, 23]. Others relied solely on data-driven methods, which, due to a lack of physical interpretation, limit generalizability to broader scenarios [15, 19]. Another study modeled underactuated finger motion with stiffness to infer contact location, but its scope was constrained to localization under specific conditions without addressing other aspects of interaction, such as object deformation [5].
@@
In this study, we present an anthropomorphic underactuated hand that achieves comprehensive situational awareness of hand–object interaction and realizes practical grasp functionalities utilizing tendon-based proprioception. To this end, we designed a miniature proprioceptive cable-driven SEA (MPC-SEA) that can be seamlessly integrated within the hand, providing direct measurement of tendon tension with high accuracy and reliability. Its compactness is achieved through a simplified sensing and guiding scheme, while ensuring stable performance under repeated operation. When combined with simple underactuated fingers without any tactile sensors or joint encoders, the MPC-SEA forms a proprioceptive SEA finger (PSF) module that provides immediate sensing capability to sensorless hands, as shown in Fig. 1.
@@
The PSF module acquires real-time tension and excursion measurements, which are combined with potential energy–based modeling of the underactuated finger to generate physically interpretable signals. These signals are further analyzed to infer grasp states of hand–object interactions, thereby enabling grasp functionalities. From a single tension–excursion profile, the hand can detect the timing of proximal and distal contacts, estimate each joint angle to reconstruct grasp posture, extract relative stiffness information, and identify external disturbances through SEA-driven compliance. All functions are executed in real time through a lightweight rule-based algorithm that combines computational simplicity with effective state estimation and stable grasp control. The proposed approach was validated through finger-level experiments—confirming finger modeling results and tension profile characterization—as well as extensive hand-level demonstrations, including grasp posture estimation of irregular objects, safe grasp force regulation for soft deformable objects without fracture, and blind grasping with proprioceptive-only recognition of objects varying in size, shape, and stiffness.
II MPC-SEA
@@ II-A MPC-SEA Design
The physical appearance and specifications of the MPC-SEA are shown in Fig. 2(a). The module measures $48.8\text{\,}\mathrm{mm}\text{/}$ $\times$ $29.0\text{\,}\mathrm{mm}\text{/}$ $\times$ $17.6\text{\,}\mathrm{mm}\text{/}$ and weighs $37.1\text{\,}\mathrm{g}\text{/}$, making it highly compact and lightweight. It can deliver up to $49\text{\,}\mathrm{N}\text{/}$ of cable tension and pull the cable at a speed of $28.9\text{\,}\mathrm{mm}\text{/}\mathrm{s}$. The structure of the MPC-SEA is shown in Fig. 2(b). Actuation is achieved by winding a cable (Dynext PU $0.9\text{\,}\mathrm{mm}\text{/}$, Amare Ropes, Italy) around a motor spool driven by a DC motor with a magnetic encoder (380:1 Micro Metal Gearmotor with 12 CPR Encoder, Pololu, USA). The cable is routed through a moving pulley housed in a sliding block, which translates linearly and compresses the spring when tension is applied, as shown in Fig. 2(c). Oilless bushings and pins guide linear motion, and the displacement of the sliding block is measured using a film-type potentiometer (ThinPot, Spectra Symbol, USA) via a spring-loaded plunger, ensuring consistent surface contact. The measured displacement is converted into cable tension using Hooke’s law, with the effective spring constant determined by four parallel springs and the two-to-one pulley mechanism, as expressed in (1).
$F_{\mathrm{tension}}$ $=\frac{F_{\mathrm{spring}}}{2}=\frac{4k_{\mathrm{sea}}\,\Delta x_{\mathrm{sb}}}{2}=2k_{\mathrm{sea}}\,\Delta x_{\mathrm{sb}}$ (1)
where $F_{\mathrm{tension}}$ is the cable tension, $F_{\mathrm{spring}}$ is the restoring force of springs, $k_{\mathrm{sea}}$ is the spring constant of a single spring, and $\Delta x_{\mathrm{sb}}$ is the displacement of the sliding block.
The linear motion is implemented using oilless bushings and guide pins instead of a conventional ball screw, and displacement sensing relies on a film-type potentiometer with a plunger. This design selection achieves significant miniaturization compared to conventional linear SEAs, while maintaining high sensing accuracy and reliability. As a self-contained module, the MPC-SEA can be used immediately by simply connecting the cable to a finger.
II-B MPC-SEA Experiment
To evaluate the accuracy and reliability of the MPC-SEA’s tension sensing, a cyclic loading test was conducted. As shown in Fig. 3(a), the cable of a fixed MPC-SEA was connected to a load cell ( [^3] FDX, KTOYO, Korea). Each cycle consisted of loading from $0\text{\,}\mathrm{N}\text{/}$ to $45\text{\,}\mathrm{N}\text{/}$ ($\approx$90% of maximum tension) and then unloading back to $0\text{\,}\mathrm{N}\text{/}$. The tension was recorded over loading, which is relevant to grasping, and compared with the reference tension measured by the load cell. A total of 50,000 cycles was performed.
III Finger Design, Modeling, and Experiment
III-A Finger and Thumb Design
@@
Unlike the other fingers, the thumb features a single degree-of-freedom interphalangeal (IP) joint for flexion and an additional carpometacarpal (CMC) rotation joint for opposition, as shown in Fig. 4(b). The CMC rotation is actuated by a worm-gear DC motor (WG12F-1510E, Motorbank, Korea). Due to spatial constraints, the thumb cannot be directly connected to the MPC-SEA; instead, it is routed through an idler pulley. In this configuration, the tendon path passes through the CMC joint axis, thereby decoupling flexion from rotation.
III-B Finger Modeling
To model the relationship between tendon excursion, tendon tension, and joint angles of the two underactuated joints, the minimum potential energy principle was applied, consistent with previous approaches in [12, 24]. Friction was neglected, and quasi-static movement was assumed. The system’s total energy $E_{\mathrm{Total}}$ comprises the elastic energy $E_{\mathrm{e}}$ from the SEA spring and the extension torsional springs, along with gravitational energy $E_{\mathrm{g}}$, as expressed in (2), (3), (4).
@@ III-B Finger Modeling
where the subscripts 1 and 2 denote the MCP and PIP joints, respectively; $\theta_{1}$ and $\theta_{2}$ are the joint angles with initial offset $\theta_{\mathrm{i1}}$ and $\theta_{\mathrm{i2}}$; $x$ is the tendon excursion; $\ell(\theta_{1},\theta_{2})$ is the tendon length required to achieve the corresponding joint angles; $k_{\mathrm{sea}}$ is the SEA spring constant; $k_{1}$ and $k_{2}$ are the torsional spring constants with preloaded angles $\theta_{\mathrm{pre1}}$ and $\theta_{\mathrm{pre2}}$; $m_{1}$ and $m_{2}$ are the phalanx masses with center-of-mass position vectors $L_{\mathrm{c1}}$ and $L_{\mathrm{c2}}$; $L_{1}$ is the position vector from the MCP to the PIP joint; $g$ is the gravitational acceleration; $R_{\mathrm{hand}}$ is the rotation matrix of the hand wrist; $R_{\mathrm{j}}(\cdot)$ is the rotation matrix about the finger joint axis; and $\mathbf{\hat{z}}$ is the unit vector along the gravitational $z$-direction.
@@
The optimization problem, formulated as shown in (5), aims to minimize the total energy of the system. The tendon excursion was discretized into small steps, and at each step, the state with the minimum energy was selected as the next step’s state. Consequently, the only decision variables were the two joint angles, since the SEA spring displacement—and thus the tension—could be readily obtained by $\Delta x_{\mathrm{sb}}=(x-\ell(\theta_{1},\theta_{2}))/2$. The constraints were defined by the joint range of motion, expressed as $\mathrm{RoM}_{j}$, and by the elastic limits of the SEA spring, expressed as $x_{\mathrm{sb,max}}$. This nonlinear optimization problem was solved in MATLAB using the sequential quadratic programming (SQP) algorithm. The modeling results are presented together with the experimental results in Section III-D.
@@
Figure 5: Tension profile characterization and grasp algorithm. (a) Tension profiles reflecting contact events for different grasp types. (b) Tension profile reflecting object deformation and relative stiffness. (c) Tension profile reflecting finger configuration changes by external disturbances. (d) Contact-pause strategy to prevent object displacement and maintain the initial location. (e) Rule-based grasping algorithm.
III-C Tension Profile Characterization
Fig. 5(a) illustrates the profiles of finger–object contact events for different grasp types. In an underactuated finger, the proximal and distal phalanges can sequentially contact the object, allowing the finger to adaptively wrap around it, or contact can occur only at the distal phalanx [25]. Contact is indicated by deviation from the free-flexion reference. Proximal contact constrains MCP motion, preventing the finger from following its original minimum-energy trajectory, which results in a slight upward shift of the tension profile. Distal contact halts finger motion, leaving only the SEA spring to compress and producing a sharp slope increase with an inflection point. The corresponding contact excursion can then be mapped to the angles of both joints with $l(\theta_{1},\theta_{2})$.

### v2

# Underactuated Robotic Hand with Grasp State Estimation Using Tendon-Based Proprioception

@@ Abstract
Anthropomorphic underactuated hands are valued for their structural simplicity and inherent adaptability. However, the uncertainty arising from interdependent joint motions makes it challenging to capture various grasp states during hand–object interaction without increasing structural complexity through multiple embedded sensors. This motivates the need for an approach that can extract rich grasp-state information from a single sensing source while preserving the simplicity of underactuation. This study proposes an anthropomorphic underactuated hand that achieves comprehensive grasp state estimation, using only tendon-based proprioception provided by series elastic actuators (SEAs). Our approach is enabled by the design of a compact SEA with high accuracy and reliability that can be seamlessly integrated into sensorless fingers. By coupling accurate proprioceptive measurements with potential energy-based modeling, the system estimates multiple key grasp state variables, including contact timing, joint angles, relative object stiffness, and external disturbances. Finger-level experimental validations and extensive hand-level grasp functionality demonstrations confirmed the effectiveness of the proposed approach. These results highlight tendon-based proprioception as a compact and robust sensing modality for practical manipulation without reliance on vision or tactile feedback.
@@ I Introduction
Anthropomorphic robotic hands have been widely adopted to replicate the functionality of the human hand. Among various actuation strategies, underactuated hands are extensively employed due to their structural simplicity and adaptability to diverse object geometries [1, 2]. As manipulation fundamentally involves physical interaction between the hand and the object, it is essential to sense grasp states throughout this interaction [3]. For instance, capturing finger configuration and force information is crucial for understanding object properties and interaction dynamics, thereby supporting the decision-making process during manipulation [4]. In underactuated hands, however, the inherent uncertainty arising from interdependent joint motions complicates grasp state sensing [1, 5]. For this reason, key grasp state variables, such as object contact, joint angles, and internal forces, are difficult to capture without embedding multiple dedicated sensors within each finger.
Yet, embedding multiple sensors into each finger makes the system more complex and less efficient, thereby undermining the fundamental advantage of underactuated mechanisms, which lies in their simple structure. This motivates sensing approaches that exploit a single comprehensive sensing signal, providing an effective alternative tailored to the characteristics of underactuation. In this context, proprioception, which provides internal information about force or position in the human muscle-tendon system [6], is a compelling sensing modality for underactuated hands over other modalities, such as vision and touch. Vision does not provide direct force feedback, making it difficult to understand hand–object interaction, and it is also vulnerable to occlusion once the hand closes around the object [7, 8]. Tactile sensors, which require additional mounting and wiring on the fingers, increase structural complexity and offer only local information rather than comprehensive information across the entire finger [9, 10].
@@
Therefore, proprioception has been widely employed in various robotic hands and grippers as a practical sensing alternative. In previous work, motor current has been used as a surrogate for proprioception, but its accuracy is constrained by electrical noise and transmission friction, which hinder reliable force sensing [11, 12]. More direct approaches–such as load cells [13] or torque sensors [5, 14], fiber Bragg gratings (FBG) sensors [10], and series elastic actuators (SEAs) [15, 16, 17, 18, 19]–have been explored for measuring internal actuation force. Among them, several studies have applied proprioceptive sensing to underactuated hands through different approaches. Some reported force profiles associated with object contact or stiffness, focusing primarily on illustrating overall trends of how force signals change during interaction [10, 17, 19]. Others relied solely on data-driven methods to infer object-related properties from proprioceptive signals, without incorporation of physical interpretation of underactuation [13, 14]. There also exist modeling-based approaches that investigated underactuated finger motion with joint stiffness to estimate contact location under specific hardware conditions [5].
However, previous studies offered only limited interpretations of an implicit proprioceptive signal that aggregates information from multiple joints. Consequently, none of these works presented a generalized estimation method that leverages the characteristics of underactuation to extract diverse grasp states from a single sensing signal. Moreover, many existing implementations result in form factors of bulky grippers or hands with enlarged bases and palms, severely limiting their applicability in anthropomorphic hands. Therefore, achieving compact in-hand integration of proprioceptive sensing while ensuring high sensing performance also remains challenging.
@@
To address the aforementioned challenges, we present an anthropomorphic underactuated robotic hand that achieves comprehensive grasp state estimation during hand–object interaction and realizes practical grasp functionalities utilizing only tendon-based proprioception. To this end, we designed a miniature proprioceptive cable-driven SEA (MPC-SEA) that can be seamlessly integrated within the hand. Its compactness is achieved through a simplified sensing and guiding scheme relative to conventional SEA structures, while ensuring tendon tension sensing with high accuracy and reliability. When attached to a simple underactuated finger without any tactile sensors or joint encoders, the MPC-SEA forms a proprioceptive SEA finger (PSF) module that enables compact in-hand integration while providing immediate sensing capability to otherwise sensorless hands, as shown in Fig. 1.
@@
The PSF module acquires real-time tendon excursion, which refers to the amount of tendon pulled by the actuator, and tendon tension. These tendon-based proprioceptive signals are combined with potential energy–based modeling to estimate physically interpretable grasp state variables from the implicit tension profile of an underactuated finger. From a single tension–excursion profile, the hand can detect the timing of proximal and distal contacts, estimate each joint angle, extract relative stiffness information, and identify external disturbances through SEA-driven compliance. All functions are executed in real time through a lightweight rule-based algorithm that combines computational simplicity with effective state estimation and stable grasp control. The proposed approach was validated through finger-level experiments—confirming finger modeling results and tension profile characterization—as well as extensive hand-level demonstrations, including grasp posture reconstruction of irregular objects, safe grasp force regulation for soft deformable objects without fracture, and blind grasping with proprioceptive-only recognition of objects varying in size, shape, and stiffness.
II Hardware Design
@@ II-A MPC-SEA Design
The physical appearance of the MPC-SEA is shown in Fig. 2(a). The module measures $48.8\text{\,}\mathrm{mm}\text{/}$ $\times$ $29.0\text{\,}\mathrm{mm}\text{/}$ $\times$ $17.6\text{\,}\mathrm{mm}\text{/}$ and weighs $37.1\text{\,}\mathrm{g}\text{/}$, making it highly compact and lightweight. It can deliver up to $49\text{\,}\mathrm{N}\text{/}$ of cable tension and pull the cable at a speed of $28.9\text{\,}\mathrm{mm}\text{/}\mathrm{s}$. The structure of the MPC-SEA is shown in Fig. 2(c). Actuation is achieved by winding a cable (Dynext PU $0.9\text{\,}\mathrm{mm}\text{/}$, Amare Ropes, Italy) around a motor spool driven by a DC motor with a magnetic encoder (380:1 Micro Metal Gearmotor with 12 CPR Encoder, Pololu, USA). The cable is routed through a moving pulley housed in a sliding block, which translates linearly and compresses the spring when tension is applied, as shown in Fig. 2(b). Oilless bushings and pins guide linear motion, and the displacement of the sliding block is measured using a film-type potentiometer (ThinPot, Spectra Symbol, USA) via a spring-loaded plunger, ensuring consistent surface contact. The measured displacement is converted into cable tension using Hooke’s law, with the effective spring constant determined by four parallel springs and the two-to-one pulley mechanism, as expressed in (1), (2), (3).
$\Delta x_{\mathrm{sb}}$ $=\frac{V_{\mathrm{o}}}{V_{\mathrm{ref}}}L_{p}-x_{0}$ (1)
$F_{\mathrm{spring}}$ $=4\times k_{\mathrm{sea}}\,\Delta x_{\mathrm{sb}}$ (2)
$T$ $=\frac{F_{\mathrm{spring}}}{2}=\frac{2k_{\mathrm{sea}}L_{p}}{V_{\mathrm{ref}}}V_{\mathrm{o}}-2k_{\mathrm{sea}}x_{0}$ (3)
where $\Delta x_{\mathrm{sb}}$ is the displacement of the sliding block, $F_{\mathrm{spring}}$ is the restoring force of springs, and $T$ is the cable tension. $V_{\mathrm{ref}}$ is the voltage reference, $V_{\mathrm{o}}$ is the voltage output, $L_{p}$ is the potentiometer length, $x_{0}$ is the initial position of the sliding block, and $k_{\mathrm{sea}}$ is the spring constant of a single spring.
The linear motion is implemented using oilless bushings and guide pins instead of a conventional ball screw, and displacement sensing relies on a film-type potentiometer with a plunger. This design selection enables significant miniaturization compared to conventional linear SEA designs, while maintaining high sensing accuracy and reliability. As a self-contained module, the MPC-SEA can be used immediately by simply connecting the cable to a finger.
II-B Finger and Thumb Design
@@
Unlike the other fingers, the thumb features a single degree-of-freedom interphalangeal (IP) joint for flexion and an additional carpometacarpal (CMC) rotation joint for opposition, as shown in Fig. 3(b). The CMC rotation is actuated by a worm-gear DC motor (603:1 WG12F-1510E, Motorbank, Korea). Due to spatial constraints, the thumb cannot be directly connected to the MPC-SEA; instead, it is routed through an idler pulley. In this configuration, the tendon path passes through the CMC joint axis, thereby decoupling flexion from rotation. The specifications of the SEA and the hand are summarized in Table I.
TABLE I: Specifications of the Hardware
Category Feature Value
Size SEA Dimension [ $\text{\,}\mathrm{mm}\text{/}$] 48.8 × 29.0 × 17.6
Finger Length (3rd) [ $\text{\,}\mathrm{mm}\text{/}$] 98.5
Hand Length [ $\text{\,}\mathrm{mm}\text{/}$] 213.5
Weight SEA Weight [ $\text{\,}\mathrm{g}\text{/}$] 37.1
Finger Weight [ $\text{\,}\mathrm{g}\text{/}$] 28.6
Hand Weight [ $\text{\,}\mathrm{g}\text{/}$] 513
Actuation SEA Max. Tension [ $\text{\,}\mathrm{N}\text{/}$] 49
SEA No-Load Speed [ $\text{\,}\mathrm{mm}\text{/}\mathrm{s}$] 28.9
Finger Closing Time [ $\text{\,}\mathrm{s}\text{/}$] 1.3
SEA Spring Constant [ $\text{\,}\mathrm{N}\text{/}\mathrm{mm}$] 5.44
Mechanism Joint Degrees of Freedom 10
Number of Actuators 6
Finger MCP / PIP Joint RoM [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 90 / 100
Finger Joint Initial Angle [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 15
Thumb IP Joint RoM [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 77
Thumb IP Joint Initial Angle [ $\text{\,}\mathrm{\SIUnitSymbolDegree}\text{/}$] 10
Finger Joint Stiffness [ $\text{\,}\mathrm{N}\text{\,}\mathrm{mm}\text{/}\mathrm{\SIUnitSymbolDegree}$] 0.86
II-C Electronics Integration
Figure 4: Electronic system architecture of the proposed hand.
Fig. 4 illustrates the overall electronic system architecture of the proposed hand. The system is powered either by an $11.1\text{\,}\mathrm{V}\text{/}$ Li-Po battery ($900\text{\,}\mathrm{m}\mathrm{A}\mathrm{h}$) or a $12\text{\,}\mathrm{V}\text{/}$ external power supply, which is regulated to $5\text{\,}\mathrm{V}\text{/}$ through a D24V10F5 regulator for all low-voltage electronics. A Teensy 4.1 serves as the main controller, executing the $1\text{\,}\mathrm{kHz}\text{/}$ sensing and motor-control loops. The controller provides the reference voltage $V_{\mathrm{ref}}$ of 3.3 $3\text{\,}\mathrm{V}\text{/}$ and receives tendon tensions from potentiometers via ADS1015 12-bit ADCs. Tendon excursion is obtained directly from the motor encoder signals embedded in the MPC-SEAs. For actuation, TB6612FNG motor drivers apply PWM-driven motor currents. This architecture enables synchronized multi-finger actuation with real-time proprioceptive feedback from every SEA module.
III Modeling and Characterization
III-A Finger Modeling
To model the relationship between tendon excursion, tendon tension, and joint angles of the two underactuated joints, the minimum potential energy principle was applied, consistent with previous approaches in [20, 21]. Friction was neglected, and quasi-static movement was assumed. Table II lists the symbols used in the finger modeling process. The system’s total energy $E_{\mathrm{Total}}$ comprises the elastic energy $E_{\mathrm{e}}$ from the SEA spring and the extension torsional springs, along with gravitational energy $E_{\mathrm{g}}$, as expressed in (4), (5), (6).
@@ III-A Finger Modeling
Figure 5: Relationship between tendon length displacement and joint angles. (a) PIP joint. (b) MCP joint.
Here, the tendon is guided by rotating bushings, resulting in a nonlinear tendon length displacement function $\ell(\theta_{1},\theta_{2})$, which can be computed from geometric configurations of the MCP and PIP joints, as illustrated in Fig. 5. Let $x_{i}$ denote the initial tendon length between two phalanges and $x_{f}$ the tendon length after the joint rotates by $\theta$. The corresponding tendon length displacement can then be obtained from the difference between $x_{i}$ and $x_{f}$. For the PIP joint, the tendon length displacement can be expressed as
$\ell_{\mathrm{pip}}(\theta)\$$=x_{i,1}-(x_{f,1}+x_{f,2})$
$=\sqrt{L_{i}^{2}-r^{2}}-\left(\sqrt{L_{f}^{2}-r^{2}}+r\beta\right).$ (7)
TABLE II: Symbol Definition for Finger Modeling
*Subscripts 1 and 2 denote the MCP and PIP joints, respectively.
Symbol Description
$\theta_{1}$, $\theta_{2}$ MCP & PIP joint angles.
$\theta_{\mathrm{i1}}$, $\theta_{\mathrm{i2}}$ Initial offsets of the MCP & PIP joint angles.
$\theta_{\mathrm{c}}$ Switching angle between two bushings in MCP joint.
$\theta_{\mathrm{pre1}}$, $\theta_{\mathrm{pre2}}$ Preloaded angles of the MCP & PIP torsional springs.
$k_{1}$, $k_{2}$ MCP & PIP torsional spring constants.
$k_{\mathrm{sea}}$ SEA spring constant.
$m_{1}$, $m_{2}$ Masses of the proximal and distal phalanges.
$L_{1}$ Position vector from MCP to PIP joint.
$L_{\mathrm{c1}}$, $L_{\mathrm{c2}}$ Center-of-mass position vectors of the phalanges.
$R_{\mathrm{hand}}$ Rotation matrix of the hand wrist.
$R_{\mathrm{j}}(\cdot)$ Rotation matrix about a finger-joint axis.
$\hat{\mathbf{z}}$ Unit vector along the gravitational $z$-direction.
$g$ Gravitational acceleration.
$T$ Tendon tension.
$x$ Tendon excursion.
$\ell(\theta_{1},\theta_{2})$ Tendon length displacement required to achieve the corresponding joint angles.
$x_{i}$, $x_{f}$ Tendon length before and after joint rotation.
$L_{i}$, $L_{f}$ Distances between the centers of the routing elements (cable knot or bushing) before and after joint rotation.
$\alpha_{i}$, $\alpha_{f}$ Angle of the tendon arc in contact with the PIP routing bushing before and after joint rotation.
$\beta$ Angle between tendon path segments between phalanges before and after joint rotation.
$r$ Radius of the routing bushing.
For the MCP joint, two bushings are located at the finger junction, and the routing path changes at a specific switching angle $\theta_{\mathrm{c}}$. Therefore, the tendon length displacement must be computed separately for the two cases: $\theta<\theta_{\mathrm{c}}$ and $\theta\geq\theta_{\mathrm{c}}$. When $\theta<\theta_{\mathrm{c}}$, the tendon length displacement is given by
$\ell_{\mathrm{mcp1}}(\theta)$ $=(x_{i,1}+x_{i,2}+x_{i,3})-(x_{f,1}+x_{f,2})$
$=\left(r\alpha_{i}+\sqrt{L_{i}^{2}-4r^{2}}+r\beta\right)-\left(r\alpha_{f}+\sqrt{L_{f}^{2}-4r^{2}}\right)$
$=\sqrt{L_{i}^{2}-4r^{2}}-\sqrt{L_{f}^{2}-4r^{2}}+r(2\beta-\theta),$ (8)
and when $\theta\geq\theta_{\mathrm{c}}$,
$\ell_{\mathrm{mcp2}}(\theta-\theta_{\mathrm{c}})$ $=(x_{i,1}+x_{i,2})-(x_{f,1}+x_{f,2}+x_{f,3})$
$=\left(r\alpha_{i}+L_{i}\right)-\left(r\alpha_{f}+L_{f}+r\beta\right)$
$=L_{i}-L_{f}-r(\theta-\theta_{\mathrm{c}}).$ (9)
Summarizing the two cases, $\ell_{\mathrm{mcp}}(\theta)$ is expressed as
$\ell_{\mathrm{mcp}}(\theta)$ $=\begin{cases}\ell_{\mathrm{mcp1}}(\theta),&\text{if }\theta<\theta_{\mathrm{c}},\\[4.0pt] \ell_{\mathrm{mcp1}}(\theta_{\mathrm{c}})+\ell_{\mathrm{mcp2}}(\theta-\theta_{\mathrm{c}}),&\text{if }\theta\geq\theta_{\mathrm{c}}.\end{cases}$ (10)
and the overall tendon length displacement $\ell(\theta_{1},\theta_{2})$ is obtained as the sum of the MCP and PIP contributions.
$\ell(\theta_{1},\theta_{2})=\ell_{\mathrm{mcp}}(\theta_{1})+\ell_{\mathrm{pip}}(\theta_{2})$ (11)
@@
The optimization problem, formulated as shown in (12), aims to minimize the total energy of the system at certain tendon excursion. The tendon excursion $x$ was discretized into small steps, and at each step, the state with the minimum energy was selected as the next step’s state. Consequently, the only decision variables were the two joint angles, since the SEA spring displacement—and thus the tension—could be readily obtained by $\Delta x_{\mathrm{sb}}=(x-\ell(\theta_{1},\theta_{2}))/2$. The constraints were defined by the joint range of motion, expressed as $\mathrm{RoM}_{j}$, and by the elastic limits of the SEA spring, expressed as $x_{\mathrm{sb,max}}$. This nonlinear optimization problem was solved in MATLAB using the sequential quadratic programming (SQP) algorithm. Consequently, the MCP and PIP joint angles, $\theta_{1}(x)$ and $\theta_{2}(x)$, as well as the corresponding tendon tension $T(x)$, can be expressed as functions of the tendon excursion $x$. The profiles of modeling results are presented together with the experimental results in Section 9.
@@
III-B Tension Profile Characterization
Figure 6: Tension profile characterization. (a) Tension profiles reflecting contact events for different grasp types. (b) Tension profile reflecting object deformation and relative stiffness. (c) Tension profile reflecting finger configuration changes by external disturbances.
Fig. 6(a) illustrates the profiles of finger–object contact events for different grasp types. In an underactuated finger, the proximal and distal phalanges can sequentially contact the object, allowing the finger to adaptively wrap around it, or contact can occur only at the distal phalanx [22]. Contact is indicated by deviation from the free-flexion reference. Proximal contact constrains MCP joint motion, preventing the finger from following its original minimum-energy trajectory, which results in a slight upward shift of the tension profile. Distal contact halts finger motion, leaving only the SEA spring to compress and producing a sharp slope increase with an inflection point. The corresponding contact excursion can then be mapped to the joint angles. Let $x_{\mathrm{pc}}$ and $x_{\mathrm{dc}}$ denote the tendon excursions at proximal and distal contact, respectively, and $T_{\mathrm{pc}}$ and $T_{\mathrm{dc}}$ the corresponding tendon tensions. After proximal contact, the subsequent tendon excursion contributes only to the PIP joint motion and SEA spring compression; therefore, tendon excursion contributing to the PIP joint motion can be calculated as
$\Delta x_{\mathrm{pip}}=x_{\mathrm{dc}}-x_{\mathrm{pc}}-\frac{T_{\mathrm{dc}}-T_{\mathrm{pc}}}{k_{\mathrm{sea}}}.$ (13)
The PIP joint angle change after proximal contact can be directly obtained from the inverse of the tendon length function $\ell_{\mathrm{pip}}^{-1}(\cdot)$ derived in (7), which was approximated by a five-order polynomial to enable fast computation of the inverse function. Since $\theta_{1}$ and $\theta_{2}$ are modeled as functions of tendon excursion, the estimated joint angles after distal contact are given by
$\begin{cases}\theta_{1,\mathrm{est}}=\theta_{1}(x_{\mathrm{pc}}),\\[3.0pt] \theta_{2,\mathrm{est}}=\theta_{2}(x_{\mathrm{pc}})+\ell_{\mathrm{pip}}^{-1}\!\left(\Delta x_{\mathrm{pip}}\right).\end{cases}$ (14)